# PreGIP: Watermarking the Pretraining of Graph Neural Networks for Deep Intellectual Property Protection

### v1

@@ 3.1 Task-Free Watermarking Framework
Theoretical Analysis. We conduct the following theoretical analysis to justify the feasibility of the proposed task-free watermarking framework.
@@
We consider a $K$-layer MLP downstream classifier $f_{C}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{c}$ with 1-Lipschitz activation function (e.g., ReLU, sigmoid, tanh) that inputs graph embeddings learned by pretrained GNN encoder $f_{E}:\mathcal{G}\rightarrow\mathbb{R}^{d}$. Let $s_{a}$ denote the margin between logit score (unnormalized score before softmax) of the predicted class and the second confident class by $f_{C}$ on a watermark graph $\mathcal{G}_{w}^{a}$. The predicted class on the paired watermark graph $\mathcal{G}_{w}^{b}$ from $f_{C}$ is guaranteed to be the same as that of $\mathcal{G}_{w}^{a}$ when:
$\|f_{E}(\mathcal{G}_{w}^{a})-f_{E}(\mathcal{G}_{w}^{b})\|_{2}<\frac{1}{2}\cdot\frac{s_{a}}{\prod_{i=1}^{K}\|\mathbf{W}_{i}\|_{2}},$ (5)
@@
Please refer to the Appendix C.1 for detailed proof. According to Theorem 3.1, the downstream classifier $f_{C}$ will predict the same label for $\mathcal{G}_{w}^{a}$ and $\mathcal{G}_{w}^{b}$ as long as their representations obtained by the watermarked encoder $f_{E}$ are close enough. Additionally, the potential adoption of the $l_{2}$-norm regularization in downstream classifier training will largely constrain the value of $\|\mathbf{W}_{i}\|_{2}$, which can ease the fulfillment of the condition specified in Theorem 3.1.
3.2 Finetuning-Resistant Watermark Injection
In this subsection, we firstly analyze how the finetuning on the watermarked GNN encoder will affect the injected watermarks. Then, we introduce the complete process of finetuning-resistant watermark injection.
Impacts of Finetuning to Watermarking. To further improve the prediction accuracy on target tasks, adversaries often finetune the stolen pretrained GNN encoder. The magnitude of modifications applied to the GNN encoder are usually limited, given that the pretrained GNN encoder already offers an effective initialization. The directions of the GNN parameter updating in finetuning can be diverse, as adversaries may apply the pretrained GNN for various tasks. Therefore, following (Bansal et al., 2022), we suppose the pretrained GNN parameters $\theta$ are updated within a spherical radius $\epsilon$ in our analysis, i.e., $\|\hat{\theta}-\theta\|_{2}\leq\epsilon$, where $\hat{\theta}$ denotes the updated parameters of the GNN encoder $f_{E}$ and $\epsilon$ should be a relatively small value. Let $s_{a}$ denote the margin between logit scores of the predicted class and the second confident class on a watermark graph $\mathcal{G}_{w}^{a}$ by the downstream classifier. Following (Cohen et al., 2019), we also assume the lower bound of $s_{a}$, denoted as $\underline{s}_{a}$ is larger than 0, which is formaluted as $s_{a}\geq\underline{s_{a}}>0$. With the above assumptions and Theorem 3.1, we have the following corollary.
@@ 3.2 Finetuning-Resistant Watermark Injection
$\small\sup_{\|\delta\|\leq\epsilon}\|f_{E}(\mathcal{G}_{w}^{a};\theta+\delta)-f_{E}(\mathcal{G}_{w}^{b};\theta+\delta)\|_{2}<\frac{1}{2}\cdot\frac{\underline{s_{a}}}{\prod_{i=1}^{K}\|\mathbf{W}_{i}\|_{2}},$ (6)
@@
The steps of deriving Corollary 3.2 are in Appendix C.2.
Finetuning-Resistant Watermarking. According to Corollary 3.2, to ensure the effectiveness of watermarking after fine-tuning for downstream tasks, we need to achieve low watermarking loss for any update $\delta,\|\delta\|_{2}\leq\epsilon$, on the model parameters $\theta$. Simultaneously, the GNN encoder parameters $\theta$ should maintain low loss for the pretraining pretext tasks to preserve the utility. To achieve the above objectives, we replace the objective function of watermark embedding in Eq.(3) with the following objective function to optimize parameters $\theta$ of the GNN encoder $f_{E}$:
$\min_{\theta}\mathcal{L}_{Pre}(\mathcal{D},\theta)+\lambda\mathcal{L}_{W}(\theta+\delta^{*})$ (7)
@@
3.3 Optimization Algorithm
$\delta_{t+1}=\delta_{t}+\alpha\nabla_{\delta}\mathcal{L}_{W}(\theta+\delta_{t}),$ (8)
$\nabla_{\theta}^{outer}$ $=\nabla_{\theta}\mathcal{L}_{Pre}(\mathcal{D},\theta)+\lambda\nabla_{\theta}\mathcal{L}_{W}(\theta+\delta_{T}).$ (9)
However, the latter term in Eq.(9) is expensive to compute as $\delta_{T}$ is a function of $\theta$. We need to unroll the training procedure of $\delta_{T}$ to calculate the gradients. Following (Nichol et al., 2018), we approximate the gradient computation by:
$\nabla_{\theta}^{outer}=\nabla_{\theta}\mathcal{L}_{Pre}(\mathcal{D},\theta)+\lambda\sum_{t=1}^{T}\nabla_{\theta}\mathcal{L}_{W}(\theta+\bar{\delta}_{t}),$ (10)
where $\bar{\delta}_{t}$ indicates gradient propagation stopping. With Eq.(10), gradient decent can be applied to update encoder parameters $\theta$. Appendix A gives the training algorithm.
@@ 3.3 Optimization Algorithm
Table 1: Comparison with baselines in protecting IP of GCL-pretrained GNNs in the semi-supervised learning setting.
Scenario
Dataset
Metrics (%)
Non-Watermarked
GCBA
Deepsign
RWM-Pre
CWM-Pre
PreGIP
Fix PROTEINS Accuracy
70.5$\pm$1.7
70.3$\pm$1.5
69.2$\pm$1.5
70.1$\pm$1.7
69.2$\pm$1.7
70.3$\pm$1.1
IP Gap
3.0$\pm$1.0
11.9$\pm$0.5
16.4$\pm$1.3
32.4$\pm$1.1
46.5$\pm$2.2
IP ROC
80.0
85.1
87.7
93.4
100
NCI1 Accuracy
66.2$\pm$0.9
65.8$\pm$0.4
65.0$\pm$0.8
65.7$\pm$1.1
65.8$\pm$0.9
65.9$\pm$0.4
IP Gap
2.2$\pm$1.6
1.6$\pm$0.2
16.6$\pm$7.9
6.0$\pm$3.9
60.0$\pm$4.9
IP ROC
56.2
59.8
85.5
64.8
100
FRANKS. Accuracy
60.8$\pm$1.3
59.6$\pm$0.7
60.5$\pm$1.2
60.0$\pm$1.0
59.9$\pm$1.1
60.5$\pm$0.9
IP Gap
4.0$\pm$2.8
13.6$\pm$1.6
16.0$\pm$2.7
5.2$\pm$3.4
79.5$\pm$24.2
IP ROC
65.9
91.4
81.9
49.8
100
Finetune PROTEINS Accuracy
70.3$\pm$1.8
67.5$\pm$1.8
69.2$\pm$1.5
69.3$\pm$2.0
69.8$\pm$1.7
69.6$\pm$0.4
IP Gap
7.3$\pm$2.9
18.4$\pm$1.7
6.6$\pm$5.7
13.0$\pm$3.7
55.5$\pm$13.5
IP ROC
94.0
85.6
58.5
71.4
94.5
NCI1 Accuracy
67.6$\pm$1.0
64.3$\pm$1.1
67.3$\pm$1.4
67.9$\pm$1.3
68.0$\pm$1.2
68.8$\pm$0.9
IP Gap
16.2$\pm$3.0
14.8$\pm$7.2
1.3$\pm$0.6
4.1$\pm$2.4
28.5$\pm$0.8
IP ROC
68.2
81.2
47.5
59.9
99.0
FRANKS. Accuracy
61.7$\pm$1.0
59.3$\pm$1.5
60.7$\pm$1.3
61.1$\pm$1.5
58.6$\pm$1.2
61.1$\pm$1.5
IP Gap
3.2$\pm$1.9
6.7$\pm$3.8
10.7$\pm$9.5
21.3$\pm$13.6
27.5$\pm$4.3
IP ROC
55.5
58.0
68.3
56.2
89.5
4 Experiments
  • Q1: Can our PreGIP accurately identify piracy models which utilize the watermarked pretrained GNN for their downstream classification tasks under various scenarios?
  • Q2: How does the number of the watermark graph pairs affect the performance of PreGIP in IP protection?
4.1 Experimental Setup
Datasets. We conduct experiments on widely used benchmark datasets for semi-supervised (You et al., 2020) and transfer learning (Hu et al., 2020) on graph classification. Specifically, for the semi-supervised setting, we use three public benchmarks, i.e., PROTEINS, NCI1 and FRANKENSTEIN (Sergey et al., 2020). For the transfer learning setting, we use 200K unlabeled molecules sampled from the ZINC15 dataset (Sterling & Irwin, 2015) for pretraining, and use 6 public molecular graph classification benchmarks (Wu et al., 2018) for the downstream tasks. More details of these datasets and learning settings are in Appendix D.1.
Baselines. As the watermarking methods on pretraining GNN encoder are rather limited, we propose several baselines adapted from existing watermarking methods. Specifically, we compare with DeepSign (Rouhani et al., 2018) which embeds watermarks into the hidden-layer activations. The verification phase of DeepSign is adjusted for black-box verification. We also adopt a sate-of-the-art graph model watermarking approach (Zhao et al., 2021). It is originally proposed for watermarking supervised classifiers which enforces random graphs to be predicted to a certain class. We extend it to a version named RWM-Pre for unsupervised encoder watermarking. To show the effectiveness of our finetuning-resistant watermark injection, we further apply the unmovable watermarking technique (Bansal et al., 2022) to RWM-Pre, resulting a baseline CWM-Pre. Furthermore, we employ the graph contrastive backdoor attack method, GCBA (Zhang et al., 2023), as a baseline for watermarking pretrained GNNs. More descriptions and implementation details of these baselines are given in Appendix D.2.
Implementation Details. We conduct experiments in watermarking GNN models pretrained by GraphCL (You et al., 2020) and edge prediction (Hu et al., 2020), which are representative works in contrastive pretraining and generative pretraining, respectively. Following GraphCL (You et al., 2020), a 2-layer GIN is employed as the GNN encoder. We configure the number of watermark graph pairs as $20$ and set $\lambda=0.1$, $\epsilon=2$ for all experiments. The hyperparameter analysis is given in Appendix E.1. To make a fair comparison, the size of watermark graphs for baselines are also set as 20. Their hyperparameters are also tuned for fair comparisons. More details can be found in Appendix D.3.
Evaluation Metrics. Firstly, we evaluate the quality of the embeddings produced by the watermarked GNN encoder. The pretraining with watermark injection is conducted on the pretraining dataset. Then, with the training set for the downstream task, we train linear classifiers that on top of the pretrained GNN encoder. The accuracy of the linear classifier on the test set for the downstream task is used to assess embeddings generated by the watermarked GNN encoder. We compute the average accuracy across 50 different train/test splits on the downstream task. Secondly, we evaluate whether the watermarking methods can distinguish the piracy models from independent models. In the evaluation, a piracy model $f_{A}$ is built by adopting watermarked pretrained GNN in downstream task classification. We obtain multiple piracy models by varying splits in downstream task. An independent classifier $f_{G}$ will use an independently pretrained GNN that is not watermarked. By varying dataset splits and random seeds in pretraining, we obtain multiple independent models. We adopts the following metrics to evaluate the watermarking performance, i.e., whether one can distinguish the piracy models from independent models:
  • IP Gap calculates the gap between the piracy model $f_{A}$ and the independent GNN $f_{G}$ in IP indication score by IP $\text{IP Gap}=|I(f_{A})-I(f_{G})|$, where $I(\cdot)\in[0,1]$ is the IP indication score of the watermarking method to infer whether the test model is copied from the protected model. The computation of $I(\cdot)$ for PreGIP and baselines is given in Eq.(4) and Appendix D.3. A larger IP Gap indicates a stronger watermarking method.
  • IP ROC evaluates how well the IP indication score $I$ can separate the piracy classifiers and independent classifiers. Specifically, the built piracy classifiers and independent classifiers are viewed as positive samples and negative samples, respectively. Then, ROC score is applied based on the IP indication score.
Table 2: Comparison with baselines in protecting IP of GCL-pretrained GNNs in the transfer learning setting.
Dataset
Metrics (%)
Non-Watermarked
Deepsign
RWM-Pre
CWM-Pre
PreGIP
Tox21 Accuracy
92.8$\pm$0.2
92.2$\pm$0.1
92.8$\pm$0.2
92.7$\pm$0.2
92.8$\pm$0.1
IP Gap
0.0$\pm$0.0
14.3$\pm$9.7
8.1$\pm$10.2
8.1$\pm$7.1
IP ROC
50.0
56.0
74.0
80.0
ToxCast Accuracy
83.4$\pm$0.2
83.2$\pm$0.2
83.5$\pm$0.2
83.3$\pm$0.1
83.5$\pm$0.2
IP Gap
1.1$\pm$0.4
4.7$\pm$4.0
3.6$\pm$0.7
7.3$\pm$3.1
IP ROC
95.0
36.0
16.0
100
BBBP Accuracy
84.9$\pm$0.8
81.9$\pm$2.2
84.7$\pm$0.9
84.3$\pm$1.3
84.5$\pm$0.6
IP Gap
0.0$\pm$0.0
51.0$\pm$36.9
25.0$\pm$11.4
49.0$\pm$4.1
IP ROC
50.0
95.0
88.0
100
BACE Accuracy
73.4$\pm$0.9
71.2$\pm$2.1
74.3$\pm$2.9
74.9$\pm$1.6
74.3$\pm$2.1
IP Gap
7.0$\pm$4.0
16.8$\pm$12.3
14.5$\pm$10.0
23.0$\pm$18.8
IP ROC
20.0
44.0
34.0
92.0
SIDER Accuracy
75.0$\pm$0.8
74.8$\pm$0.2
75.0$\pm$0.4
74.8$\pm$0.2
75.0$\pm$0.4
IP Gap
1.4$\pm$0.6
2.1$\pm$1.6
4.3$\pm$3.3
18.3$\pm$10.5
IP ROC
68.0
48.0
64.0
88.0
ClinTox Accuracy
92.3$\pm$1.2
91.9$\pm$1.3
91.7$\pm$1.0
92.1$\pm$1.3
92.2$\pm$1.3
IP Gap
0.0$\pm$0.0
21.0$\pm$7.0
11.5$\pm$13.2
37.5$\pm$43.4
IP ROC
50.0
48.0
80.0
80.0
4.2 Results of Watermarking Pretraining
To answer Q1, we first compare PreGIP with baselines in watermarking the contrastive pretraining for the semi-supervised graph classification. Then, we conduct experiments in watermarking generative pretraining to show the flexibility of PreGIP. Additionally, pretraining for the transfer learning is also evaluated to demonstrate the transferability of PreGIP in watermarking.
Pretraining for Semi-Supervised Learning. Following (You et al., 2020), a scheme of semi-supervised learning for downstream graph classification is implemented. Particularly, the GNN is firstly pretrained on the whole dataset $\mathcal{D}$ without using labels. Then, the pretrained GNN is adopted to the downstream classifier. Only a small set of samples $\mathcal{D}_{L}\subset\mathcal{D}$ with downstream task labels are available for the training of the classifier. During the training of the classifier for downstream tasks, the parameters of pretrained GNN encoder can be fixed or finetuned. Both scenarios of Fix and Finetune are considered in our experiments. More details about the semi-supervised setting are in Appendix D.3. 50 piracy and independent models are build to compute metrics as described in Sec. 4.1. The results of watermarking GraphCL are reported in Tab. 1. We also report the test accuracy of pretrained GNNs without watermarking as a reference. From the table, we observe that:
  • [leftmargin=*]
  • In both Fix and Finetune scenarios, PreGIP achieves accuracy similar to the pretrained GNN without watermarking. This demonstrates that PreGIP can obtain watermarked GNN encoder with high quality embeddings to facilitate the downstream tasks.
(a) IP ROC
(b) Accuracy
Figure 3: Watermarking generative pretraining on PROTEINS.
Flexibility in Pretraining Strategies. To show the flexibility of PreGIP to different pretraining strategies, we conduct experiments on a representative generative pretraining method, i.e., edge prediction (Hu et al., 2020). Semi-supervised graph classification is also applied as the downstream task. All the other settings are kept the same as Appendix D.3. Note that GCBA is designed for graph contrastive learning and not feasible to the edge prediction pretraining. We report IP ROC and accuracy on PROTEINS in Fig. 3. More results on other datasets can be found in Appendix E.2. From the figure, we observe that PreGIP consistently achieves better IP ROC and accuracy compared to baselines in both Fix and Finetune scenarios. It demonstrates the effectiveness of PreGIP for the edge prediction pretraining scheme, which implies the flexibility of PreGIP to various pretraining strategies.
Pretraining for Transfer Learning. We also evaluate the performance of PreGIP in the transfer learning setting of molecular property prediction (You et al., 2020). In the transfer learning setting, we first pretrain the GNN in ZINC15 dataset that contains 200k unlabeled molecules, and then finetune the model in 6 molecular datasets (i.e., Tox21, ToxCast, BBBP, BACE, SIDER and ClinTox). GraphCL is set as the pretraining strategy. 10 privacy classifiers and independent classifiers are built to compute the IP Gap and IP ROC. All other evaluation settings are the same as Sec. 4.1. More details of the transfer learning setting are in Appendix D.1. The results on the 6 datasets are reported in Tab. 2. Compared with the baselines, the pretrained GNN encoder watermarked by PreGIP can effectively identify piracy models even the watermarked encoder is updated under the transfer learning setting. Meanwhile, performance on the downstream tasks are comparable with non-watermarked pretrained GNN. This validates the watermarking effectiveness of PreGIP in the transfer learning setting.
4.3 Impact of the Number of Watermark Graph Pairs
To answer Q2, we conduct experiments to explore the watermarking performance of PreGIP given different budgets in the number of watermark graph pairs. Specifically, we vary the number of watermarking graphs as $\{1,5,10,20,50,100\}$. We set GraphCL as the pretraining scheme and the semi-supervised graph classification as the downstream task. The other settings are described in Sec. 4.1. Fig. 4 reports the IP Gap and accuracy on PROTEINS and FRANKENSTEIN under semi-supervised setting in the finetuning scenario. More results of IP ROC is given in Appendix E.4. As the number of watermarking graphs increases, the IP Gap consistently rises, while accuracy slightly decreases with more watermark graph pairs in training. This suggests more watermark graph pairs can strengthen the watermarking, but too many may negatively affect the embedding space of the pretrained GNN. Hence, an appropriate number of watermark graph pairs is necessary to balance the IP protection and model utility.
(a) IP Gap
(b) Accuracy
Figure 4: Impact of the number of watermark graph pairs.
4.4 Ablation Studies
[…]
  • [leftmargin=*]
[…]
(a) IP ROC
(b) Accuracy
Figure 5: Ablation study results on PROTEINS.
5 Related Works
Graph Model Pretraining. Graph model pretraining (You et al., 2020; Sun et al., 2023a; Xia et al., 2022) aims to learn useful knowledge from large-scale data to facilitate various downstream tasks. During the pretraining phase, a pretext task such as graph construction will be deployed to train GNN encoder for a high quality embedding space. Based on the types of pretext tasks, graph pretraining methods can be categorized into contrastive pretraining and generative pretraining. Graph contrastive learning methods (You et al., 2020; Qiu et al., 2020) aim to maximize the mutual information between different augmentation views of graphs. For instance, GraphCL (You et al., 2020) testifies the pretraining performance with different combinations of augmentation views. Recently, All in One (Sun et al., 2023a) designs a multi-task prompting framework to unify node-level, edge-level, and graph level tasks in a pretrained contrastive learning model. The generative pretraining aims to train graph model by generative tasks such as graph reconstruction and attribute prediction. For example, edge prediction and masked node prediction are investigated in (Hu et al., 2020). GraphMAE (Hou et al., 2022) deploys a task of feature reconstruction with re-mask decoding strategy. As for the introduction about the related works in graph neural networks, it can be found in the Appendix B.
Model Watermarking. Because of the great values of deep learning models, various model watermarking methods (Uchida et al., 2017; Zhang et al., 2018; Adi et al., 2018; Rouhani et al., 2018; Sun et al., 2023b) are designed for deep IP protection. Watermarking neural networks are firstly developed by Uchida et al., which regularizes the model parameters to inject the watermark. Then, during the verification phase, an extractor is applied to the model parameters to verify the existence of the watermark. To achieve black-box verification that does not require the model parameters of suspect models, backdoor-based watermark methods (Zhang et al., 2018; Adi et al., 2018) are proposed. Specifically, they enforce trigger samples to be predicted as a predefined class. Predictions on trigger samples are used as IP message to verify the model ownership. To resist watermark removal by adversaries, random noises are added to model parameters during the watermark embedding (Bansal et al., 2022). However, this method is not applicable to the black-box verification on self-supervised pretraining. Recently, SSLGuard (Cong et al., 2022) introduces a watermarking scheme for self-supervised learning pretrained encoders.
The aforementioned frameworks focus on independent and identically distributed data. They cannot be directly applicable graph structured data because of the significant differences in input samples and network architectures. Recently, some initial efforts employ watermarking to protect the IP of supervised GNN classifier (Xu et al., 2023a; Zhao et al., 2021). Following backdoor-based methods on images (Adi et al., 2018),  Xu et al. enforce GNN classifier to predict trigger graphs as certain class to achieve watermarking. Our PreGIP is inherently different from the aforementioned methods: (i) we focus on watermarking the pretraining of GNN encoder without the downstream classification task information; (ii) we propose a novel finetuning-resistant watermark injection framework, which can identify unauthorized adoption of pretraining GNN encoder even after finetuning on downstream tasks.
6 Conclusion and Future Work
In this paper, we study a novel problem of watermarking the pretraining of GNNs for deep IP protection. Specifically, we propose a task-free model verification framework that can watermark the GNN pretraining without the downstream task information. A watermarking loss is applied to embed the watermarks without degrading the quality of representations obtained by the GNN encoder. Moreover, a finetuning-resistant watermarking approach is implemented to guarantee the identification of any suspect model derived from finetuning the protected GNN encoder across various tasks. Extensive experiments on different datasets and learning settings demonstrate that PreGIP can effectively watermark the pretraining of GNNs. In the future, it would be promising to investigate potential watermark removal strategies to bypass the identification of pirated models.
7 Broader Impact
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Training Algorithm of PreGIP
@@
0:  Pretraining dataset $\mathcal{D}$, pretraining method, hyperparameters $\lambda$ and $\epsilon$
@@
8:     Obtain the approximate gradients for GNN encoder $f_{E}$ by Eq.(10)

### v2

@@ 3.1.4. Ownership Verification
3.1.5. Theoretical Analysis
We conduct the following theoretical analysis to justify the feasibility of the proposed task-free watermarking framework.
@@ 3.1.5. Theoretical Analysis
We consider a $K$-layer MLP downstream classifier $f_{C}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{c}$ with 1-Lipschitz activation function (e.g., ReLU, sigmoid, tanh) that inputs graph embeddings learned by pretrained GNN encoder $f_{E}:\mathcal{G}\rightarrow\mathbb{R}^{d}$. Let $s_{a}$ denote the margin between the logit score (unnormalized score before softmax) of the predicted class and the second confident class by $f_{C}$ on a watermark graph $\mathcal{G}_{w}^{a}$. The predicted class on the paired watermark graph $\mathcal{G}_{w}^{b}$ from $f_{C}$ is guaranteed to be the same as that of $\mathcal{G}_{w}^{a}$ when:
(5) $\small\|f_{E}(\mathcal{G}_{w}^{a})-f_{E}(\mathcal{G}_{w}^{b})\|_{2}<\frac{1}{2}\cdot\frac{s_{a}}{\prod_{i=1}^{K}\|\mathbf{W}_{i}\|_{2}},$
@@
Please refer to the Appendix A.1 for detailed proof. According to Theorem 3.1, the downstream classifier $f_{C}$ will predict the same label for $\mathcal{G}_{w}^{a}$ and $\mathcal{G}_{w}^{b}$ as long as their representations obtained by the watermarked encoder $f_{E}$ are close enough. Additionally, the potential adoption of the $l_{2}$-norm regularization in downstream classifier training will largely constrain the value of $\|\mathbf{W}_{i}\|_{2}$, which can ease the fulfillment of the condition specified in Theorem 3.1.
3.2. Finetuning-Resistant Watermark Injection
In this subsection, we first analyze how the finetuning on the watermarked GNN encoder will affect the injected watermarks. Then, we introduce the finetuning-resistant watermark injection.
3.2.1. Impacts of Finetuning to Watermarking
To further improve the prediction accuracy on target tasks, adversaries often finetune the stolen pretrained GNN encoder. The magnitude of modifications applied to the GNN encoder is usually limited, given that the pretrained GNN encoder already offers an effective initialization. The directions of the GNN parameter updating in finetuning can be diverse, as adversaries may apply the pretrained GNN for various tasks. Therefore, following (Bansal et al., 2022), we suppose the pretrained GNN parameters $\theta$ are updated within a spherical radius $\epsilon$ in our analysis, i.e., $\|\hat{\theta}-\theta\|_{2}\leq\epsilon$, where $\hat{\theta}$ denotes the updated parameters of the GNN encoder $f_{E}$ and $\epsilon$ should be a relatively small value. Let $s_{a}$ denote the margin between logit scores of the predicted class and the second confident class on a watermark graph $\mathcal{G}_{w}^{a}$ by the downstream classifier. Following (Cohen et al., 2019), we also assume the lower bound of $s_{a}$, denoted as $\underline{s}_{a}$ is larger than 0, which is formaluted as $s_{a}\geq\underline{s_{a}}>0$. With the above assumptions and Theorem 3.1, we have the following corollary.
@@ 3.2.1. Impacts of Finetuning to Watermarking
(6) $\small\sup_{\|\delta\|\leq\epsilon}\|f_{E}(\mathcal{G}_{w}^{a};\theta+\delta)-f_{E}(\mathcal{G}_{w}^{b};\theta+\delta)\|_{2}<\frac{1}{2}\cdot\frac{\underline{s_{a}}}{\prod_{i=1}^{K}\|\mathbf{W}_{i}\|_{2}},$
@@
The steps of deriving Corollary 3.2 are in Appendix A.2.
3.2.2. Finetuning-Resistant Watermarking
According to Corollary 3.2, to ensure the effectiveness of watermarking after fine-tuning for downstream tasks, we need to achieve low watermarking loss for any update $\delta,\|\delta\|_{2}\leq\epsilon$, on the model parameters $\theta$. Simultaneously, the GNN encoder parameters $\theta$ should maintain low loss for the pretraining pretext tasks to preserve the utility. To achieve the above objectives, we replace the objective function of watermark embedding in Eq.(3) with the following objective function to optimize parameters $\theta$ of the GNN encoder $f_{E}$:
(7) $\min_{\theta}\mathcal{L}_{Pre}(\mathcal{D},\theta)+\lambda\mathcal{L}_{W}(\theta+\delta^{*})$
@@ 3.2.2. Finetuning-Resistant Watermarking
Table 1. Results in protecting IP of GCL-pretrained GNNs under the semi-supervised learning setting.
Scenario Dataset Metrics (%) Non-Watermarked WPE SSL-WM GCBA Deepsign RWM-Pre CWM-Pre PreGIP
Fix PROTEINS Accuracy 70.5$\pm$1.7 69.7$\pm$1.6 59.7$\pm$2.0 70.3$\pm$1.5 69.2$\pm$1.5 70.1$\pm$1.7 69.2$\pm$1.7 70.3$\pm$1.1
IP Gap - 18.4$\pm$16.2 23.6$\pm$3.0 3.0$\pm$1.0 11.9$\pm$0.5 16.4$\pm$1.3 32.4$\pm$1.1 46.5$\pm$2.2
IP ROC - 79.1 100 80.0 85.1 87.7 93.4 100
NCI1 Accuracy 66.2$\pm$1.1 64.1 $\pm$1.9 63.3 $\pm$3.3 65.8$\pm$0.4 65.0$\pm$0.8 65.7$\pm$1.1 65.8$\pm$0.9 65.9$\pm$0.4
IP Gap - 14.5$\pm$5.6 2.3$\pm$6.2 2.2$\pm$1.6 1.6$\pm$0.2 16.6$\pm$7.9 6.0$\pm$3.9 60.0$\pm$4.9
IP ROC - 62.1 69.1 56.2 59.8 85.5 64.8 100
FRANKS. Accuracy 60.8$\pm$1.3 59.5$\pm$1.1 60.3$\pm$0.9 59.6$\pm$0.7 60.5$\pm$1.2 60.0$\pm$1.0 59.9$\pm$1.1 60.5$\pm$0.9
IP Gap - 25.9$\pm$20.8 4.5$\pm$1.5 4.0$\pm$2.8 13.6$\pm$1.6 16.0$\pm$2.7 5.2$\pm$3.4 79.5$\pm$24.2
IP ROC - 69.8 40.3 65.9 91.4 81.9 49.8 100
Finetune PROTEINS Accuracy 70.3$\pm$1.8 64.6$\pm$1.2 64.7$\pm$1.2 67.5$\pm$1.8 69.2$\pm$1.5 69.3$\pm$2.0 69.8$\pm$1.7 69.6$\pm$0.4
IP Gap - 15.3$\pm$8.8 2.6$\pm$2.7 7.3$\pm$2.9 18.4$\pm$1.7 6.6$\pm$5.7 13.0$\pm$3.7 55.5$\pm$13.5
IP ROC - 88.8 68.2 94.0 85.6 58.5 71.4 94.5
NCI1 Accuracy 67.6$\pm$1.0 67.5$\pm$1.3 67.7$\pm$1.3 64.3$\pm$1.1 67.3$\pm$1.4 67.9$\pm$1.3 68.0$\pm$1.2 68.8$\pm$0.9
IP Gap - 8.5$\pm$7.4 8.4$\pm$5.6 16.2$\pm$3.0 14.8$\pm$7.2 1.3$\pm$0.6 4.1$\pm$2.4 28.5$\pm$0.8
IP ROC - 51.8 52.6 68.2 81.2 52.5 59.9 99.0
FRANKS. Accuracy 61.7$\pm$1.0 59.6$\pm$1.6 61.5$\pm$1.4 59.3$\pm$1.5 60.7$\pm$1.3 61.1$\pm$1.5 58.6$\pm$1.2 61.1$\pm$1.5
IP Gap - 25.2$\pm$21.1 26.3$\pm$25.9 3.2$\pm$1.9 6.7$\pm$3.8 10.7$\pm$9.5 21.3$\pm$13.6 27.5$\pm$4.3
IP ROC - 61.8 83.0 55.5 58.0 68.3 56.2 89.5
@@
0:  Pretraining dataset $\mathcal{D}$, pretraining method, $\lambda$, $\epsilon$
@@
8:     Obtain the gradients for GNN encoder $f_{E}$ by Eq.(10)